two seperate question please answer both either abcd
Find the angle between x = 0 and y= O 71/6 O 72 O 77/4 O 7/3
Find a unit vector in the direction of the vector • ฟ- - เก | เก
Answer #1

12 11 [:] у [:] nl = √1270²+12 = √2 151 = √0²+12+1 = 2 Coso = 20- fullg/ 11 OX1 + 0*1 + 1x1 ปx VE Losil = ta 2 0 : 7/3 let اک